# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ I.2 Near-Constant Weight Design
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{v\in[n_{2}]\setminus\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{[}\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\big{(}v-n_{2}\rho\big{)}(1+o(1))\Big{]}:=\mathcal{T}_{11}-\mathcal{T}_{12}.$ (430)
Note that $\mathcal{T}_{12}$ is again insignificant, since analogously to (420) and (428) we have
$|\mathcal{T}_{12}|\leq O(n)\mathbb{P}(\mathscr{A}_{2}^{c})\leq O(n)\exp\Big{(}-\frac{C_{*,2}\ell n}{k}\Big{)}=o\Big{(}\frac{\ell n}{k}\Big{)}.$ (431)
Observe that $\mathcal{T}_{11}$ can be written as an expectation:
$\mathcal{T}_{11}$ $=\mathbbm{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}\Big{[}\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\big{(}V-n_{2}\rho\big{)}(1+o(1))\Big{]}$ (432)
$=\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\big{(}\mathbbm{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}(V)-n_{2}\rho\big{)}(1+o(1)).$ (433)
Thus, it suffices to show $\mathbbm{E}[\mathcal{T}_{11}]$ is lower bounded by the right-hand side of (347), and now we calculate $\mathbb{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}[V]$. While $V$ represents the number of $1$s in the noisy outcomes $\mathbf{Y}_{2}$, it is useful to also define $U$ as the number of tests in $\mathbf{Y}_{2}$ that would be positive in the absence of noise (i.e., the tests corresponding to $\mathbf{X}_{s_{\mathrm{eq}}}=0,\mathbf{X}_{s_{\mathrm{dif}}}\neq 0$). Then, conditioned on $U=u$, the effect of noise gives
$\mathbb{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},U=u}[V]$ $=\mathbb{E}\big{[}\mathrm{Bin}(u,1-\rho)+\mathrm{Bin}(n_{2}-u,\rho)\big{]}$ (434)
$=u(1-\rho)+(n_{2}-u)\rho$ (435)
$=(1-2\rho)u+\rho n_{2}.$ (436)
@@
$\mathbb{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}[V]=(1-2\rho)\mathbbm{E}_{U|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}[U]+\rho n_{2}.$ (437)
@@
$\geq\sum_{i=1}^{\ell\Delta}\frac{n_{2}-i+1}{n}$ (439)
$=\frac{n_{2}\ell\Delta}{n}-\frac{(\ell\Delta-1)\ell\Delta}{2n}$ (440)
$\geq\frac{\ell\nu n_{2}}{k}\Big{(}1-\frac{\ell\nu n}{2kn_{2}}\Big{)},$ (441)
where (439) holds since at most $i-1$ out of the $n_{2}$ tests are ruled out by the previous $i-1$ placements. Substituting (437) and (441) into (433) gives
$\mathcal{T}_{11}$ $\geq(1+o(1))\log\Big{(}\frac{1-\rho}{\rho}\Big{)}(1-2\rho)\frac{\ell\nu n_{2}}{k}\Big{(}1-\frac{\ell\nu n}{2kn_{2}}\Big{)}$ (442)
$=\left(\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log\frac{1-\rho}{\rho}\right)\big{(}1+o(1)\big{)},$ (443)
where (443) holds because we are considering $n_{2}=ne^{-\nu}(1+o(1))$ (as explained earlier in (420)).
We have now established that the right-hand side of (443) is a lower bound for $I_{\ell}^{n}$ (with other terms such as $\mathcal{T}_{12}$ being factored into the $o(1)$ part). This matches the upper bound on $I_{\ell}^{n}$ derived above, and the proof is complete.
J.3 Proof of Lemma 13 (Characterization of $\mathbb{P}(V)$)
We specialize some of the developments from the proof of Lemma 11 to the case that $s_{\mathrm{eq}}=\varnothing$, $s_{\mathrm{dif}}=s$, and $\ell=k$. Hence, in the notation used therein, we have $\mathbf{Y}_{1}=\varnothing$ and $\mathbf{Y}_{2}=\mathbf{Y}$, $M$ denotes the number of tests that would be positive if there were no noise, and $V$ represents for the number of positive tests in the noisy results (i.e., the number of 1s in $\mathbf{Y}$). Our goal is to show that $\mathbb{P}(v)=e^{-o(n)}$ for all $v$ within a high-probability set (i.e., $V$ lies in that set with $1-o(1)$ probability). Since we trivially have $\mathbb{P}(v)\leq 1$, it suffices to only lower bound $\mathbb{P}(v)$.
By setting $\delta=n^{-1/4}=o(1)$ in (359), we have with $1-o(1)$ probability that
$M=(1-e^{-\nu}+o(1))n.$ (444)
Moreover, by setting $\delta=n^{-1/4}=o(1)$ in (363), we have with $1-o(1)$ probability that
$V=(\rho\star e^{-\nu}+o(1))n,$ (445)
@@
Give $M=m$, we have that $V$ follows a sum of two independent binomial variables: $(V|M=m)=B_{1}+B_{2}$, where $(B_{1}|M=m)\sim\mathrm{Bin}(m,1-\rho)$ and $(B_{2}|M=m)\sim\mathrm{Bin}(n-m,\rho)$. Given fixed $m$ and $v$ values satisfying (444)–(445), we can find $b_{1},b_{2}$ with $b_{1}+b_{2}=v$ such that
$b_{1}=n(1-\rho)(1-e^{-\nu})(1+o(1)),\quad b_{2}=n\rho e^{-\nu}(1+o(1)),$ (446)
i.e., such that the two are asymptotically close to the mean values of $\mathrm{Bin}(m,1-\rho)$ and $\mathrm{Bin}(n-m,\rho)$, and also sum to $v$. Letting $\mathcal{M}_{\rm typical}$ be a high-probability set of $m$ values that all satisfy (444), we can then lower bound $\mathbb{P}(V=v)$ as follows:
$\mathbb{P}(V=v)$ $\geq\sum_{m\in\mathcal{M}_{\rm typical}}\mathbb{P}(M=m)\mathbb{P}(B_{1}=b_{1}|M=m)\mathbb{P}(B_{2}=b_{2}|M=m).$ (447)
We now observe that $\mathbb{P}(B_{1}=b_{1}|M=m)$ and $\mathbb{P}(B_{2}=b_{2}|M=m)$ both scale as $e^{-o(n)}$; this directly follows from the anti-concentration of binomial random variables (Lemma 1) and $b_{1},b_{2}$ being arbitrarily close to the respective means (see (446)). Since $\mathcal{M}_{\rm typical}$ is a high-probability set of $m$ values, we deduce from (447) that $\mathbb{P}(V=v)\geq e^{-o(n)}$, which completes the proof.
Acknowledgment
This work was supported by the Singapore National Research Foundation (NRF) under grant number A-0008064-00-00.

### 23 Oct 2024 v2

@@ I.2 Near-Constant Weight Design
$\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ -\sum_{v\in[n_{2}]\setminus\mathcal{V}_{\mathbf{x}_{s_{\mathrm{eq}}}}}\mathbb{P}(V=v)\Big{[}\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\big{(}v-n_{2}\rho\big{)}(1+o(1))\Big{]}:=\mathcal{T}_{11}-\mathcal{T}_{12}.$ (J.79)
Note that $\mathcal{T}_{12}$ is again insignificant, since analogously to (J.69) and (J.77) we have
$|\mathcal{T}_{12}|\leq O(n)\mathbb{P}(\mathscr{A}_{2}^{c})\leq O(n)\exp\Big{(}-\frac{C_{*,2}\ell n}{k}\Big{)}=o\Big{(}\frac{\ell n}{k}\Big{)}.$ (J.80)
For $\mathcal{T}_{11}$, we write
$\mathcal{T}_{11}$ $=\mathbbm{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}\Big{[}\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\big{(}V-n_{2}\rho\big{)}(1+o(1))\Big{]}$ (J.81)
$=\log\Big{(}\frac{1-\rho}{\rho}\Big{)}\big{(}\mathbbm{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}(V)-n_{2}\rho\big{)}(1+o(1)).$ (J.82)
Thus, it suffices to show $\mathbbm{E}[\mathcal{T}_{11}]$ is lower bounded by the right-hand side of (H.8), and now we calculate $\mathbb{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}[V]$. While $V$ represents the number of $1$s in the noisy outcomes $\mathbf{Y}_{2}$, it is useful to also define $U$ as the number of tests in $\mathbf{Y}_{2}$ that would be positive in the absence of noise (i.e., the tests corresponding to $\mathbf{X}_{s_{\mathrm{eq}}}=0,\mathbf{X}_{s_{\mathrm{dif}}}\neq 0$). Then, conditioned on $U=u$, the effect of noise gives
$\mathbb{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}},U=u}[V]$ $=\mathbb{E}\big{[}\mathrm{Bin}(u,1-\rho)+\mathrm{Bin}(n_{2}-u,\rho)\big{]}$ (J.83)
$=u(1-\rho)+(n_{2}-u)\rho$ (J.84)
$=(1-2\rho)u+\rho n_{2}.$ (J.85)
@@
$\mathbb{E}_{V|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}[V]=(1-2\rho)\mathbbm{E}_{U|\mathbf{X}_{s_{\mathrm{eq}}}=\mathbf{x}_{s_{\mathrm{eq}}}}[U]+\rho n_{2}.$ (J.86)
@@
$\geq\sum_{i=1}^{\ell\Delta}\frac{n_{2}-i+1}{n}$ (J.88)
$=\frac{n_{2}\ell\Delta}{n}-\frac{(\ell\Delta-1)\ell\Delta}{2n}$ (J.89)
$\geq\frac{\ell\nu n_{2}}{k}\Big{(}1-\frac{\ell\nu n}{2kn_{2}}\Big{)},$ (J.90)
where (J.88) holds since at most $i-1$ out of the $n_{2}$ tests are ruled out by the previous $i-1$ placements, and (J.90) uses $\Delta=\frac{\nu n}{k}$. Substituting (J.86) and (J.90) into (J.82) gives
$\mathcal{T}_{11}$ $\geq(1+o(1))\log\Big{(}\frac{1-\rho}{\rho}\Big{)}(1-2\rho)\frac{\ell\nu n_{2}}{k}\Big{(}1-\frac{\ell\nu n}{2kn_{2}}\Big{)}$ (J.91)
$=\left(\frac{n\nu e^{-\nu}\ell}{k}(1-2\rho)\log\frac{1-\rho}{\rho}\right)\big{(}1+o(1)\big{)},$ (J.92)
where (J.92) holds because we are considering $n_{2}=ne^{-\nu}(1+o(1))$ (as explained earlier in (J.69)).
We have now established that the right-hand side of (J.92) is a lower bound for $I_{\ell}^{n}$ (with other terms such as $\mathcal{T}_{12}$ being factored into the $o(1)$ part). This matches the upper bound on $I_{\ell}^{n}$ derived above, and the proof is complete.
J.3 Proof of Lemma I.2 (Characterization of $\mathbb{P}(V)$)
We specialize some of the developments from the proof of Lemma H.3 to the case that $s_{\mathrm{eq}}=\varnothing$, $s_{\mathrm{dif}}=s$, and $\ell=k$. Hence, in the notation used therein, we have $\mathbf{Y}_{1}=\varnothing$ and $\mathbf{Y}_{2}=\mathbf{Y}$, $M$ denotes the number of tests that would be positive if there were no noise, and $V$ represents for the number of positive tests in the noisy results (i.e., the number of 1s in $\mathbf{Y}$). Our goal is to show that $\mathbb{P}(v)=e^{-o(n)}$ for all $v$ within a high-probability set (i.e., $V$ lies in that set with $1-o(1)$ probability). Since we trivially have $\mathbb{P}(v)\leq 1$, it suffices to only lower bound $\mathbb{P}(v)$.
By setting $\delta=n^{-1/4}=o(1)$ in (J.8), we have with $1-o(1)$ probability that
$M=(1-e^{-\nu}+o(1))n.$ (J.93)
Moreover, by setting $\delta=n^{-1/4}=o(1)$ in (J.12), we have with $1-o(1)$ probability that
$V=(\rho\star e^{-\nu}+o(1))n,$ (J.94)
@@
Give $M=m$, we have that $V$ follows a sum of two independent binomial variables: $(V|M=m)=B_{1}+B_{2}$, where $(B_{1}|M=m)\sim\mathrm{Bin}(m,1-\rho)$ and $(B_{2}|M=m)\sim\mathrm{Bin}(n-m,\rho)$. Given fixed $m$ and $v$ values satisfying (J.93)–(J.94), we can find $b_{1},b_{2}$ with $b_{1}+b_{2}=v$ such that
$b_{1}=n(1-\rho)(1-e^{-\nu})(1+o(1)),\quad b_{2}=n\rho e^{-\nu}(1+o(1)),$ (J.95)
i.e., such that the two are asymptotically close to the mean values of $\mathrm{Bin}(m,1-\rho)$ and $\mathrm{Bin}(n-m,\rho)$, and also sum to $v$. Letting $\mathcal{M}_{\rm typical}$ be a high-probability set of $m$ values that all satisfy (J.93), we can then lower bound $\mathbb{P}(V=v)$ as follows:
$\mathbb{P}(V=v)$ $\geq\sum_{m\in\mathcal{M}_{\rm typical}}\mathbb{P}(M=m)\mathbb{P}(B_{1}=b_{1}|M=m)\mathbb{P}(B_{2}=b_{2}|M=m).$ (J.96)
We now observe that $\mathbb{P}(B_{1}=b_{1}|M=m)$ and $\mathbb{P}(B_{2}=b_{2}|M=m)$ both scale as $e^{-o(n)}$; this directly follows from the anti-concentration of binomial random variables (Lemma A.1) and $b_{1},b_{2}$ being arbitrarily close to the respective means (see (J.95)). Since $\mathcal{M}_{\rm typical}$ is a high-probability set of $m$ values, we deduce from (J.96) that $\mathbb{P}(V=v)\geq e^{-o(n)}$, which completes the proof.